A Bibliography of Publications of Jan Flusser

Jan Flusser

Tel: ?n/a? FAX: ?n/a?

E-mail: flusser@utia.cas.cz (Internet)

19 February 2013 Version 1.05

Abstract

This bibliography records publications of Jan Flusser.

Title word cross-reference

1-D [FS96].

Activity [KSvF91]. Adaptive [Flu92a]. Affine [FS93b, FS93a, FSv93, FS93c, FS94a, FS94b, FSS94a]. analysis [FS98a]. application [KŠF98]. Approach [Flu91b, FS94b, FS98a]. archaeological [HF97]. atmospheric [KvF97]. Aurorae [KSvF91].

Based [Flu91b, FS94b, FSS94b, Flu98a, SF96b]. best [FS97b, FS98b]. between [SF96a]. binary [Flu98b]. blind [KvF97, KŠF98]. blur [FSS95a, FSS95b, FSS96b]. blurred [FSS94b, FSS96a, KZFP98]. boundary [Flu98a]. boundary-based [Flu98a]. calculation [Flu98a, Flu98b]. case [FS97b]. Character [FS93a, FS94a]. Classification [FS93b, FS97a]. Coding [MF93b]. coefficients [Flu95]. Complex [Flu91a]. control [SF96a]. correspondence [SF96a]. corrupted [KvF97].

D [FS96]. deconvolution [KŠF98]. decreasing [FS97b]. deformed [SF95a, SF96a, SF96b]. Degraded [FS98a, FS96, FSS96b, FS97a]. Description [Flu92b, FSS94b]. Detection [ZKPF99, KZFP98]. digital [FS96]. direct [HF98]. distortion [FS94b, FSS94a]. Distortions [Flu91a].

Effective [Flu98a]. Effects [KSvF91]. ellipses [HF98]. environment [FS98b]. Estimation [HF97]. Experimental [FSv93].

Fast [Flu91a, Flu98b]. Feature [Flu91b, KZFP98]. Feature-Based [Flu91b]. features [FSS94b, FSS95a, FS97b, FS98b, SF95a, SF96b].

Finite [MF93a, MF93b]. fitting [HF98]. Flare [KSvF91].

Geometric [Flu91a, FS94b, FSS94a, Flu98b].

Half [KvF97]. Half-blind [KvF97].

Image [Flu91b, Flu92a, FSS95a, MF93a, FS98a]. Images [Flu91a, ZKPF99, FS94b, FSS94b, FSS94a, FSS96a, FSS96b, Flu98b, KZFP98, KvF97, KŠF98, SF96a]. Invariant [Flu92b, FSS95a, FS98a]. Invariants [FS93b, FS93a, FSv93, FS93c, FSS95b, FS96, FS94a, FSS94a, FS97a, SF95b, SF97].

least [HF98]. likelihood [Flu95]. linear [FSS96b].

Matching [Flu91b, Flu95]. means [FSS94a, Flu95]. Measure [Flu92b]. Method [Flu92a, FSS96a, FS97a, KŠF98]. Minimum [KSvF91]. Moment [FS93b, FS93a, FSv93, FS93c, FSS94b, FS94a, FS94b, FSS94a]. Moment-based [FSS94b, FS94b]. moments [FSS96a, Flu98a, Flu98b]. motion [FSS96b]. multichannel [KŠF98]. Multiframe [ZKPF99].

noisy [FS98b]. Numerically [HF98].

Object [Flu92b, Flu95, Flu98a]. Objects [FS93b].

Pattern [FS93c]. Patterns [FS93b]. permutation [SF97]. Phase [KSvF91]. Picture [MF93b]. Point [SF97, KZFP98, SF95a]. Points [ZKPF99, SF96a]. polygons [SF95b, SF96b]. pottery [HF97]. profiles [HF97]. projective [SF95b, SF97]. projectively [SF95a, SF96a, SF96b]. Proton [KSvF91].

Radon [MF93a, MF93b].

Recognition [FS93a, FS93c, FSS96a, FSS96b, FS94a, FSS94b, FS96, SF95a, SF96b]. Registration [Flu91a, Flu92a, FSS94a, FS94b, SF96a]. Representations [MF93a]. respect [FSS95a, FSS95b]. restoration [FSS96b, KvF97]. Robust [ZKPF99]. robustness [FS97b].

Search [SF96a]. Selecting [FS97b, FS98b]. sets [SF95a]. Shape [Flu92b]. sherds [HF97]. signals [FS96, FS97a]. Significant [ZKPF99]. Similarity [Flu92b]. Solar [KSvF91, KŠF98]. squares [HF98]. stable [HF98]. Study [FSv93]. symmetric [FSS95b].

Terrestrial [KSvF91]. tool [FS94a].

Transform [MF93a, MF93b].

turbulence [KvF97].

Uniqueness [FSv93].

Vertex [SF96b]. Vertex-based [SF96b]. via [MF93a].

without [FSS96b].

References

Flusser:1991:FRI

[Flu91a] Jan Flusser. A fast registration of images with complex geometric distortions. In Peter Johansen and Søren Olsen, editors, Proceedings of 7th Scandinavian Conference on Image Analysis: Aalborg, August 13–16, 1991 (SCIA ’91), pages 1011–1019. Pattern Recognition Society of Denmark, Aalborg, Denmark, August 1991.

Flusser:1991:FBA

[Flu91b] Jan Flusser. A feature-based approach to image matching. In R. Klette, editor, Computer Analysis of Images and Patterns: Proceedings of the 4th International Conference CAIP ’91, pages 227– 232. Akademie Verlag, Berlin, Germany, September 1991.

Flusser:1992:AMI

[Flu92a] Jan Flusser. An adaptive method for image registration. Pattern Recognition, 25(1):45–54, 1992. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Flusser:1992:ISD

[Flu92b] Jan Flusser. Invariant shape description and measure of object similarity. In 4th International Conference on Image Processing and its Applications, pages 139– 142. IEE, London, UK, April 1992.

Flusser:1995:OMM

[Flu95] Jan Flusser. Object matching by means of matching likelihood coefficients. Pattern Recognition Letters, 16:893–900, 1995. CODEN PRLEDG. ISSN 0167-8655 (print), 1872-7344 (electronic).

Flusser:1998:EBB

[Flu98a] J. Flusser. Effective boundarybased calculation of object moments. In R. Klette, G. Gimel’farb, and R. Kakarala, editors, Proc. Image and Vision Computing, pages 369–374. University of Auckland, Auckland, New Zealand, November 1998.

Flusser:1998:FCG

[Flu98b] J. Flusser. Fast calculation of geometric moments of binary images. In M. Gengler, editor, Pattern Recognition and Medical Computer Vision, pages 265–274. ÖCG, Illmitz, 1998.

Flusser:1993:CRA

[FS93a] Jan Flusser and Tomáš Suk. Character recognition by affine moment invariants. In D. Chetverikov and W. Kropatsch, editors, Computer Analysis of Images and Patterns, pages 572–577. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1993.

Flusser:1993:COP

[FS93b] Jan Flusser and Tomáš Suk. Classification of objects and patterns by affine moment invariants. In G. Vernazza, editor, Image Processing: Theory and Applications, pages 345–348. Elsevier Science Publishers, Amsterdam, The Netherlands, 1993.

Flusser:1993:PRA

[FS93c] Jan Flusser and Tomáš Suk. Pattern recognition by affine moment invariants. Pattern Recognition, 26 (1):167–174, 1993. CODEN PT-NRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Flusser:1994:AMI

[FS94a] Jan Flusser and Tomáš Suk. Affine moment invariants: a new tool for character recognition. Pattern Recognition Letters, 15:433– 436, 1994. CODEN PRLEDG.

ISSN 0167-8655 (print), 1872-7344 (electronic).

Flusser:1994:MBA

[FS94b] Jan Flusser and Tomáš Suk. A moment-based approach to registration of images with affine geometric distortion. IEEE Transactions on Geoscience and Remote Sensing, 32(2):382–387, 1994. CO-DEN IGRSD2. ISSN 0196-2892.

Flusser:1996:IRD

[FS96] Jan Flusser and Tomáš Suk. Invariants for recognition of degraded 1-D digital signals. In International Conference on Pattern Recognition. Proceedings, pages 389–393. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1996.

Flusser:1997:CDS

[FS97a] Jan Flusser and Tomáš Suk. Classification of degraded signals by the method of invariants. Signal Processing, 60(2):243–249, 1997. CO-DEN SPRODR. ISSN 0165-1684.

Flusser:1997:SBF

[FS97b] Jan Flusser and Tomáš Suk. Selecting the best features in the case of decreasing robustness. In P. Pudil, J. Novovičová, and J. Grim, editors, Proc. 1st IAPR TC1 Workshop on Statistical Techniques in Pattern Recognition, pages 49–54. ÚTIA AV ČR, Praha, Czech Republic, June 1997.

Flusser:1998:DIA

[FS98a] J. Flusser and T. Suk. Degraded image analysis: An invariant approach. IEEE Transactions on Pattern Analysis and Machine Intelligence, 20(6):590–603, 1998. CO-DEN ITPIDJ. ISSN 0162-8828.

Flusser:1998:SBF

[FS98b] J. Flusser and T. Suk. On selecting the best features in a noisy environment. Kybernetika, 34(4):441–446, 1998. CODEN KYBNAI. ISSN 0023-5954.

Flusser:1994:RIA

[FSS94a] Jan Flusser, Stanislav Saic, and Tomáš Suk. Registration of images with affine geometric distortion by means of moment invariants. In J. Desachy, editor, Image and Signal Processing for Remote Sensing, pages 843–852. SPIE Optical Engineering Press, Bellingham, WA, USA, September 1994.

Flusser:1994:MBF

[FSS94b] Jan Flusser, Tomáš Suk, and Stanislav Saic. Moment-based features for description and recognition of blurred images. In R. Klette, W. Kropatsek, and F. Solina, editors, Theoretical Foundations of Computer Vision, page 5. Dagstuhl, March 1994.

Flusser:1995:IFI

[FSS95a] Jan Flusser, Tomáš Suk, and Stanislav Saic. Image features invariant with respect to blur. Pattern Recognition, 28(11):1723– 1732, 1995. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Flusser:1995:IRS

[FSS95b] Jan Flusser, Tomáš Suk, and Stanislav Saic. Invariants with respect to symmetric blur. In G. Borgefors, editor, Proceedings of the 9th Scandinavian Conference on Image Analysis, pages 341– 348. SSAB, Uppsala, Sweden, June 1995.

Flusser:1996:RBI

[FSS96a] Jan Flusser, Tomáš Suk, and Stanislav Saic. Recognition of blurred images by the method of moments. IEEE Transactions on Image Processing, 5(3):533–538, 1996. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Flusser:1996:RID

[FSS96b] Jan Flusser, Tomáš Suk, and Stanislav Saic. Recognition of images degraded by linear motion blur without restoration. Computing, 11:37–51, 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Flusser:1993:ESU

[FSv93] Jan Flusser, Tomáš Suk, and Stanislava Šimberová. An experimental study on uniqueness of the affine moment invariants. In V. Hlaváč and T. Pajdla, editors, Czech Pattern Recognition Workshop ’93, pages 119–125. Czechoslovak Pattern Recognition Society, Praha, Czech Republic, November 1993.

Halir:1997:EPS

[HF97] R. Halíř and Jan Flusser. Estimation of profiles of sherds of archaeological pottery. In T. Pajdla, editor, Proc. Czech Pattern Recognition Workshop ’97, pages 126–130. Czech Pattern Recognition Society, Praha, Czech Republic, February 1997.

Halir:1998:NSD

[HF98] R. Halíř and J. Flusser. Numerically stable direct least squares fitting of ellipses. In Proc. 6th International Conference in Central Europe on Computer Graphics and Visualization. WSCG ’98, pages 125–132. CZ, Plzeň, Czech Republic, February 1998.

Krejci:1998:NMB

[KŠF98] R. Krejčí, S. Šimberová, and J. Flusser. A new multichannel blind deconvolution method and its application to solar images. In A. Jain, S. Venkatesh, and B. C. Lovell, editors, Proc. 14th Intl. Conf. Pattern Recognition, pages 1765–1767. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1998.

Krivsky:1991:PFT

[KSvF91] L. Křivský, M. A. Soliman, Stanislava Šimberová, and Jan Flusser. Proton flare and terrestrial effects with aurorae in the minimum phase of solar activity. Bulletin of the Astronomical Institutes of Czechoslovakia, 42(6):386–389, 1991.

Krejci:1997:HBR

[KvF97] Radovan Krejčí, Stanislava Šimberová, and Jan Flusser. Half-blind restoration of images corrupted by atmospheric turbulence. In V. Krasnoproshin, J. Soldek, S. Ablameyko, and V. Shmerko, editors, Pattern Recognition and Information Processing, pages 384– 387. WUPS, Szczecin, Poland, May 1997.

Kautsky:1998:FPD

[KZFP98] J. Kautsky, B. Zitová, J. Flusser, and G. Peters. Feature point detection in blurred images. In R. Klette, G. Gimel’farb, and R. Kakarala, editors, Proc. Image and Vision Computing, pages 103–108. University of Auckland, Auckland, New Zealand, November 1998.

Matus:1993:IRF

[MF93a] František Matúš and Jan Flusser.

Image representations via a finite radon transform. IEEE Transactions on Pattern Analysis and Machine Intelligence, 15(10):996– 1006, 1993. CODEN ITPIDJ. ISSN 0162-8828.

Matus:1993:PCF

[MF93b] František Matúš and Jan Flusser.

Picture coding by a finite radon transform. In General Assembly of the International Union of Radio Science, page 106. URSI, Kyoto, Japan, August 1993.

Suk:1995:FRP

[SF95a] Tomáš Suk and Jan Flusser. The features for recognition of projectively deformed point sets. In B. Werner, editor, Proc. IEEE Intl. Conf. Image Processing, pages 348– 351. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1995.

Suk:1995:PIP

[SF95b] Tomáš Suk and Jan Flusser. The projective invariants for polygons. In V. Hlaváč and R. Šára, editors, Computer Analysis of Images and Patterns, pages 729–734. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1995.

Suk:1996:SCBa

[SF96a] Tomáš Suk and Jan Flusser. Search of the correspondence between the control points for registration of the projectively deformed images. In J. Desachy, editor, Image and Signal Processing for Remote Sensing, pages 28–37. SPIE Optical Engineering Press, Bellingham, WA, USA, September 1996.

Suk:1996:SCBb

[SF96b] Tomáš Suk and Jan Flusser. Vertex-based features for recognition of projectively deformed polygons. Pattern Recognition, 29(3): 361–367, 1996. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Suk:1997:PPP

[SF97] Tomáš Suk and Jan Flusser. Point projective and permutation invariants. In G. Sommer, K. Danilidis, and J. Pauli, editors, Computer Analysis of Images and Patterns, pages 74–81. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1997.

Zitova:1999:RDS

[ZKPF99] B. Zitová, J. Kautsky, G. Peters, and J. Flusser. Robust detection of significant points in multiframe images. Pattern Recognition Letters, 20:199–206, 1999. CODEN PRLEDG. ISSN 0167-8655 (print), 1872-7344 (electronic).